# Robust DNN Partitioning and Resource Allocation Under Uncertain Inference Time

### v1

@@ Abstract
Index Terms:
Edge intelligence, DNN partitioning, uncertain inference time, chance-constrained programming, convex-concave procedure.
@@ I Introduction
In practical systems, the inference time of DNNs is variable and uncertain, and it cannot be determined until the inference task is executed [16], [17]. In [16], the authors assess the inference time of convolutional neural networks on the SoCs, observing significant performance variations under inference time outliers. In [17], the authors observe that the inference time of various DNN models applied to autonomous driving are uncertain, and analyze several factors that influence the fluctuations in DNN inference time. Different from the object detection tasks in [17], we verify the variations in inference time of several DNN models for the classification task on the CIFAR-10 dataset using the CPU and GPU platforms, as shown in Fig. 1. We also find that the uncertainty of DNN inference time is affected by the model structure, I/O speed, hardware platform, etc. Moreover, it can be observed from Fig. 1 that the inference time of different models on different hardware exhibits significant randomness, which makes its distribution knowledge difficult to obtain accurately. Indeed, uncertain inference time brings a significant challenge to edge-device collaboration. It is well known that deciding DNN model partitioning based on worst-case inference time tends to be overly conservative, and extending task deadlines can compromise the timeliness of the system. Therefore, it is necessary to consider the robust DNN partitioning and resource allocation to provide performance guarantees.
@@ II-A DNN Model Partitioning
Extensive research focuses on DNN model partitioning and resource allocation in collaboration inference. Various works are investigated from different perspectives, such as collaborative paradigms, DNN model structures, and inference approaches. The cloud-end collaboration paradigm partially shifts DNN inference from the device to the cloud [7], [18]. In [7], the neurosurgeon algorithm is proposed to find an intermediate partitioning point in the DNN model, keeping the front-end model on the device and offloading the back-end model to the cloud. Leveraging a similar principle, [18] proposes a distributed partitioning strategy that divides the DNN model into the cloud, the edge server, and the end devices. To reduce the latency of cloud inference, the edge-end collaboration paradigm is studied [5], [15]. Edgent [5] utilizes mobile edge computing (MEC) for DNN collaborative inference through device-edge synergy by adaptive partitioning and right-sizing DNNs to reduce latency. Based on [5], [15] proposes a learning-based method that optimizes DNN partitioning, early exit point selection, and computing resource allocation.
@@ III-A DNN Model Partitioning
Different DNNs exhibit a range of structures. For example, AlexNet and VGG are organized as single chains [23], [25], while ResNet features two asymmetric branches [24]. Typically, the structure of DNNs is modeled as DAGs [14], [26]. However, this DAG-based modeling can be quite complex. For simplicity, we use the block-based modeling approach [10], [12]. This method involves dividing the DAG into multiple blocks, effectively transforming it into a serial chain structure. As shown in Fig. 4, each block we construct consists of multiple layers, including convolutional layers (Conv), pooling layers (Pool), batch normalization layers (BN), activation layers (such as ReLU), etc. Denote $M$ as the number of blocks in the DNN model. Then, the set of partitioning points is represented as $\mathcal{M}\triangleq\{0,1,\ldots,M\}$. Let $x_{n,m}\in\{0,1\},n\in\mathcal{N},m\in\mathcal{M}$ be the partitioning decision, and there is only one partitioning point for each mobile device, i.e., $\sum_{m\in\mathcal{M}}x_{n,m}=1,\forall n\in\mathcal{N}$. Specifically, $x_{n,m}=1$ indicates that mobile device $n$ executes partitioning at the $m$th point, and $x_{n,m}=0$ otherwise. For instance, $x_{n,0}=1$ means that mobile device $n$ only executes edge inference, $x_{n,M}=1$ means that mobile device $n$ only executes local inference, and $x_{n,m}=1$ means that the first $m$ blocks execute local inference, and the remaining $(M-m)$ blocks execute edge inference.
@@ III-B Inference Time and Energy Consumption
where $t_{n,0}^{\mathrm{loc}}=u_{n,0}^{\mathrm{loc}}=0$. The dynamic power consumption of the COMS circuit is denoted as $\alpha cV^{2}f$, where $\alpha$ is the activity factor, $c$ is the load capacitance, $V$ is the supply voltage, and $f$ is the CPU/GPU clock frequency [27]. Moreover, $V$ is approximately linear to the frequency when the CPU/GPU operates in the non-low frequency range, i.e., $V=kf$ [28]. Thus, the corresponding energy consumption of mobile device $n$ to execute local inference is
@@
The MEC server can generate a virtual machine (VM) for each mobile device. Each VM is configured with the corresponding DNN model to its associated mobile device and executes the offloading task in parallel. We assume each mobile device is assigned a dedicated VM, and each VM exclusively serves its corresponding mobile device. Let $u_{n,k}^{\mathrm{vm}}$ denote the edge inference time of mobile device $n$ in the $k$th block. The edge inference time of mobile device $n$ can be expressed as
@@
where $t_{n,M}^{\mathrm{vm}}=0$. The size of the inference result (e.g., the object name of the image recognition output) is often much smaller than the raw data and feature data, so the time taken to download the inference result from the edge node to mobile devices can be ignored. In addition, since the MEC server is powered by the grid, the energy consumption of edge inference and result downloading is not considered [29], [30].
@@ III-C Problem Formulation
[…]
@@ IV-A Mean Inference Time
Figure 6: The mean inference time at different partitioning points on Jeston Xavier NX CPU/GPU. Discrete points represent the experiment data, while the continuous lines represent the fitted functions.
Therefore, we utilize real-world data to model the functional relationship between inference time and CPU/GPU frequency. The inference time of widely used DNNs (i.e., AlexNet and ResNet152) is tested on multiple devices (e.g., Jeston Xavier NX CPU and GPU) by frequency scaling. Specifically, AlexNet and ResNet152 are partitioned into $2$ and $3$ blocks, respectively. The sets of partitioning points are defined as $m\in M\triangleq\{0,1,2\}$ for AlexNet and $m\in M\triangleq\{0,1,2,3\}$ for ResNet152, where $m=0$ indicates that the inference is executed on the VM, resulting in the inference time of the device being 00.111Due to space limitations, only the cases with 2 and 3 blocks are presented here. However, in the experiments where the blocks are partitioned into 8 or 9, each block demonstrated a similar curve, as depicted in Fig. 6. We use nonlinear least square method to fit the measured data above. Fig. 6 illustrates the fitting curve and coefficient of AlexNet and ResNet152 for different partitioning points on CPU and GPU. For AlexNet, the squared 2-norm of the residual at $m=1$ and $m=2$ is 2.0e-4 $\text{s}^{2}$ and 9.7e-4 $\text{s}^{2}$, respectively. For ResNet152, the squared 2-norm of the residual at $m=1$, $m=2$, and $m=3$ is 5.7e-4 $\text{s}^{2}$, 8.0e-4 $\text{s}^{2}$, and 2.9e-3 $\text{s}^{2}$, respectively.
@@ IV-B Variance and Covariance of Inference Time
To solve the above problem, we use the maximum value in the CPU/GPU frequency scaling range as the variance of the inference time. The variance of inference time when mobile device $n$ selects partitioning point $m$ is obtain by
$v_{n,m}^{\mathrm{loc}}=\max_{\forall f_{n}\in\mathcal{F}}\left\{v_{n,m}^{\mathrm{loc}}(f_{n})\right\},\forall n\in\mathcal{N},m\in\mathcal{M},$ (11)
where $v_{n,m}^{\text{loc }}(\cdot)=\mathbb{E}\left[\left(t_{n,m}^{\text{loc }}(\cdot)-\bar{t}_{n,m}^{\text{loc }}(\cdot)\right)^{2}\right]$ and $\mathcal{F}\triangleq\left[f_{\min},f_{\max}\right]$. This approximation may introduce some errors; however, simulation results show the error is acceptable. The experiments and analysis are discussed in Section VI. Note that in this work, we assume the CPU/GPU frequencies of mobile devices can be scaled, whereas the CPU/GPU frequencies of VMs remain constant. Therefore, $\overline{t}_{n,m}^{\mathrm{vm}}$ and $v_{n,m}^{\mathrm{vm}}$ can be obtained through simple online measurement.
During collaborative inference, covariance information between the mobile device and the VM is also required. Thus, we designate Jetson Xavier NX as the mobile device and RTX4080 as the VM, and calculate the covariance at different partitioning points. The experimental results show that the covariance curve closely matches the variance curve in Fig. 7. It is because the computing capacity of the VM is higher than mobile devices, leading to lower inference time and fluctuations. Therefore, similar to variance, the covariance of inference time at different partitioning points is approximated by
$w_{n,m,m^{\prime}}=\max_{\forall f_{n}\in\mathcal{F}}\left\{w_{n,m,m^{\prime}}(f_{n})\right\},\forall n\in\mathcal{N},m,m^{\prime}\in\mathcal{M}.$ (12)
@@ V-B Resource Allocation Subproblem
[…]
@@ V-C DNN Model Partitioning Subproblem
[…]
@@
where $\rho>0$ is a penalty parameter. Then, the concave terms of the constraints (35c), (35d), and (35e) are linearized to obtain convex constraints for a minimization problem and to solve a sequence of convex problems successively. Specifically, at $i$th iteration, update $\left\{\mathbf{x}^{(i)},\mathbf{y}^{(i)}\right\}$ by solving the following approximate problem, which is parameterized by $\left\{\mathbf{x}^{(i-1)},\mathbf{y}^{(i-1)}\right\}$ obtained at $(i-1)$th iteration.
@@
where $\rho^{(i-1)}$ is the penalty parameter at the $(i-1)$th iteration. Problem (36) is a convex problem that can be solved efficiently by an IPT algorithm. The pseudo-code for solving problem (36) is presented in Algorithm 1. The computational complexity of solving problem (36) using an IPT algorithm is $\mathcal{O}\left(N^{3}M^{3}\right)$, and the number of iterations of the IPT algorithm is $\mathcal{O}(\sqrt{NM}\log(1/\xi))$, where $\xi$ is the convergence accuracy. Therefore, the total computational complexity of Algorithm 1 is $\mathcal{O}\left((NM)^{3.5}\log(1/\xi)\right)$ [39]. Note that the sequence solution $\left\{\mathbf{x}^{(i)}\right\}_{i=1}^{\infty}$ to problem (36) can converge to a stationary point of problem (33), as shown in [40]. Since problem (33) and problem (24) are equivalent, Algorithm 1 can also converge to a stationary point of problem (24).
@@ VI-A Simulation Setup
TABLE III: The parameters of AlexNet on Jetson Xavier NX CPU.
@@
TABLE IV: The parameters of ResNet152 on Jetson Xavier NX GPU.
@@
Two widely-used DNNs, AlexNet [23] and ResNet152 [24], are considered. The two DNNs are fully deployed on mobile devices and the MEC server. The task of the mobile device is image recognition, which is extracted from the object recognition dataset CIFAR-10 [42]. The processing unit of mobile devices adopts Jetson Xavier NX CPU and GPU [43]. We assume that AlexNet is deployed on the Jetson Xavier NX CPU, while ResNet152 is deployed on Jetson Xavier NX GPU. The VM assigned to each mobile device uses the GeForce RTX 4080. Specific configurations are shown in Table II.
The energy efficiency coefficient $\kappa_{n}$ of Jetson Xavier NX CPU and GPU is evaluated using the power testing tool Tegrastats of NVIDIA [44]. Specifically, Jetson Xavier NX is first set to a fixed power consumption mode. Then, the power consumption of the CPU and GPU at different frequencies is measured, and finally, $\kappa_{n}$ is obtained based on the measured data. By estimation, the average $\kappa_{n}$ of the Jetson Xavier NX CPU and GPU are $0.8\times 10^{-27}\mathrm{W}/(\text{cycle}/\mathrm{sec})^{3}$ and $2.8\times 10^{-27}\mathrm{W}/(\text{cycle}/\mathrm{sec})^{3}$, respectively.
AlexNet and ResNet152 are divided into 8 and 9 blocks, corresponding to 9 and 10 partition points, respectively. The feature data size of each block can be calculated based on its output data shape. The mean inference time for each block is obtained through 500 experiments, and then the variance and covariance can also be calculated based on the mean and measured data. The specific parameters are shown in Table III and IV. Unless otherwise specified, the above parameters are used by default. For comparison, we consider the following two policies as benchmark:
@@
    Worst-case policy: the upper bound of $t_{n,m}^{\mathrm{loc}}$ and $t_{n,m}^{\mathrm{vm}}$ obtained by the experiment is taken as the inference time, and the task deadline is not allowed to be violated.
@@
    Optimal policy: the DNN partitioning is obtained using the exhaustive search method, which can find the optimal partitioning point, but its computational complexity is exponential.

### v2

@@ I Introduction
In practical systems, the inference time of DNNs is variable and uncertain, and it cannot be determined until the inference task is executed [16], [17]. In [16], the authors assess the inference time of convolutional neural networks on the SoCs, observing significant performance variations under inference time outliers. In [17], the authors observe that the inference time of various DNN models applied to autonomous driving are uncertain, and analyze several factors that influence the fluctuations in DNN inference time. Different from the object detection tasks in [17], we verify the variations in inference time of several DNN models for the classification task on the CIFAR-10 dataset using the CPU and GPU platforms, as shown in Fig. 1. We also find that the uncertainty of DNN inference time is affected by the model structure, I/O speed, hardware platform, etc. Moreover, it can be observed from Fig. 1 that the inference time of different models on different hardware exhibits significant randomness, which makes its distribution function difficult to obtain accurately. Indeed, uncertain inference time brings a significant challenge to edge-device collaboration. It is well known that deciding DNN model partitioning based on worst-case inference time tends to be overly conservative, and extending task deadlines can compromise the timeliness of the system. Therefore, it is necessary to consider the robust DNN partitioning and resource allocation to provide performance guarantees.
@@ II-A DNN Model Partitioning
Extensive research focuses on DNN model partitioning and resource allocation in collaboration inference. Various works are investigated from different perspectives, such as collaborative paradigms, DNN model structures, and inference approaches. The cloud-end collaboration paradigm partially shifts DNN inference from the device to the cloud [7], [18]. In [7], the neurosurgeon algorithm is proposed to find an intermediate partitioning point in the DNN model, keeping the front-end model on the device and offloading the back-end model to the cloud. Leveraging a similar principle, [18] proposes a distributed partitioning strategy that divides the DNN model into the cloud, the edge server, and the end devices. To reduce the latency of cloud inference, the edge-end collaboration paradigm is studied [5], [15]. Edgent [5] utilizes mobile edge computing (MEC) for DNN collaborative inference through device-edge synergy by adaptive partitioning and right-sizing DNNs to reduce latency. Based on [5], [15] proposes a learning-based method that optimizes DNN partitioning, early exit point selection, and computing resource allocation. [45] proposes an integrated sensing, computing, and communication (ISCC) architecture that jointly optimizes AI model partitioning, as well as integrated sensing and communication, to deliver low-latency intelligent services at the network edge.
@@ III-A DNN Model Partitioning
Different DNNs exhibit a range of structures. For example, AlexNet and VGG are organized as single chains [23], [25], while ResNet features two asymmetric branches [24]. Typically, the structure of DNNs is modeled as DAGs [14], [26]. However, this DAG-based modeling can be quite complex. For simplicity, we use the block-based modeling approach [10], [12]. This method involves dividing the DAG into multiple blocks, effectively transforming it into a serial chain structure. As shown in Fig. 4, each block we construct consists of multiple layers, including convolutional layers (Conv), pooling layers (Pool), batch normalization layers (BN), activation layers (such as ReLU), etc. Denote $M$ as the number of blocks in the DNN model. Then, the set of partitioning points is represented as $\mathcal{M}\triangleq\{0,1,\ldots,M\}$. Let $x_{n,m}\in\{0,1\},n\in\mathcal{N},m\in\mathcal{M}$ be the partitioning decision, and there is only one partitioning point for each mobile device, i.e., $\sum_{m\in\mathcal{M}}x_{n,m}=1,\forall n\in\mathcal{N}$. Specifically, $x_{n,m}=1$ indicates that mobile device $n$ executes partitioning at the $m$th point, and $x_{n,m}=0$ otherwise. For instance, $x_{n,0}=1$ means that mobile device $n$ only executes edge inference, $x_{n,M}=1$ means that mobile device $n$ only executes local inference, and $x_{n,m}=1$ means that the first $m$ blocks execute local inference, and the remaining (Mm)(M-m) blocks execute edge inference.
@@ III-B Inference Time and Energy Consumption
where $t_{n,0}^{\mathrm{loc}}=u_{n,0}^{\mathrm{loc}}=0$. The dynamic power consumption of the COMS circuit is denoted as $\alpha cV^{2}f$, where $\alpha$ is the activity factor, $c$ is the load capacitance, $V$ is the supply voltage, and $f$ is the CPU/GPU clock frequency [27]. Moreover, $V$ is approximately linear to the frequency when the CPU/GPU operates in the non-low frequency range, i.e., $V=kf$ [27], [28]. Thus, the corresponding energy consumption of mobile device $n$ to execute local inference is
@@
The MEC server can generate a virtual machine (VM) for each mobile device that executes edge inference. Each VM is configured with the corresponding DNN model to its associated mobile device and performs the offloading task in parallel. For simplicity, we assume that the total computing resources on the edge server are equally allocated to each VM, and each VM is dedicated to serving its corresponding mobile device [46], [47]. The delay introduced by VM configuration is usually small (e.g., $\sim 10\mathrm{~ms}$), so its impact on the overall delay is negligible and has been disregarded. [^1] For example, advanced VM configuration techniques can make its latency less than 10 ms, and techniques such as hot start and lightweight configuration can further reduce the latency of VM configuration [48], [49]. Let $u_{n,k}^{\mathrm{vm}}$ denote the edge inference time of mobile device $n$ in the $k$th block. The edge inference time of mobile device $n$ can be expressed as
@@
where $t_{n,M}^{\mathrm{vm}}=0$. The time spent downloading the task computation results from the edge node (e.g., 5G Base Station) to the device is negligible, due to the strong transmission power and high bandwidth of the edge node, as well as the relatively small output data size (compared to the raw or feature data size). For example, in object recognition applications, the output typically consists of vehicles, pedestrians, etc., which requires only a few bytes and is much smaller than the corresponding input raw data or feature data, the latter of which may be several megabytes. In addition, since the MEC server is powered by the grid, the energy consumption of edge inference and result downloading is not considered [29], [30].
@@ III-C Problem Formulation
$\ \text{s.t.}\ \,\mathbb{P}\left\{T_{n}\leq D_{n}\right\}\geq 1-\varepsilon_{n},\forall n\in\mathcal{N},$ (9b)
@@ IV-A Mean Inference Time
Figure 6: The mean inference time at different partitioning points on Jeston Xavier NX CPU/GPU. Discrete points represent the sample mean, while the continuous lines represent the fitted functions.
Therefore, we utilize real-world data to model the functional relationship between inference time and CPU/GPU frequency. The inference time of widely used DNNs (i.e., AlexNet and ResNet152) is tested on multiple devices (e.g., Jeston Xavier NX CPU and GPU) by frequency scaling. Specifically, AlexNet and ResNet152 are partitioned into $2$ and $3$ blocks, respectively. The sets of partitioning points are defined as $m\in M\triangleq\{0,1,2\}$ for AlexNet and $m\in M\triangleq\{0,1,2,3\}$ for ResNet152, where $m=0$ indicates that the inference is executed on the VM, resulting in the inference time of the device being 0.222Due to space limitations, only the cases with 2 and 3 blocks are presented here. However, in the experiments where the blocks are partitioned into 8 or 9, each block demonstrated a similar curve, as depicted in Fig. 6. We use nonlinear least square method to fit the measured data above. Fig. 6 illustrates the fitting curve and coefficient of AlexNet and ResNet152 for different partitioning points on CPU and GPU. For AlexNet, the squared 2-norm of the residual at $m=1$ and $m=2$ is 2.0e-4 $\text{s}^{2}$ and 9.7e-4 $\text{s}^{2}$, respectively. For ResNet152, the squared 2-norm of the residual at $m=1$, $m=2$, and $m=3$ is 5.7e-4 $\text{s}^{2}$, 8.0e-4 $\text{s}^{2}$, and 2.9e-3 $\text{s}^{2}$, respectively.
@@ IV-B Variance and Covariance of Inference Time
To solve the above problem, we use the mean value in the CPU/GPU frequency scaling range as the variance of the inference time. The variance of inference time when mobile device $n$ selects partitioning point $m$ is obtain by
$v_{n,m}^{\mathrm{loc}}=\frac{1}{|\mathcal{F}|}\sum_{\forall f_{n}\in\mathcal{F}}v_{n,m}^{\text{loc}}\left(f_{n}\right),\forall n\in\mathcal{N},m\in\mathcal{M},$ (11)
where $v_{n,m}^{\text{loc}}(\cdot)=\mathbb{E}\left[\left(t_{n,m}^{\text{loc}}(\cdot)-\bar{t}_{n,m}^{\text{loc}}(\cdot)\right)^{2}\right]$ and $\mathcal{F}$ is the set of measurable frequency points of the CPU/GPU. This approximation may introduce some errors; however, simulation results show the error is acceptable. The experiments and analysis are discussed in Section VI. Note that in this work, we assume the CPU/GPU frequencies of mobile devices can be scaled, whereas the CPU/GPU frequencies of VMs remain constant. Therefore, $\overline{t}_{n,m}^{\mathrm{vm}}$ and $v_{n,m}^{\mathrm{vm}}$ can be obtained through simple online measurement.
During collaborative inference, covariance information between the mobile device and the VM is also required. Thus, we designate Jetson Xavier and Nano as the mobile device and RTX4080 as the VM, and calculate the covariance at different partitioning points. The experimental results show that the covariance curve closely matches the variance curve in Fig. 7. It is because the computing capacity of the VM is higher than mobile devices, leading to lower inference time and fluctuations. Therefore, similar to variance, the covariance of inference time at different partitioning points is approximated by
$w_{n,m,m^{\prime}}=\frac{1}{|\mathcal{F}|}\sum_{\forall f_{n}\in\mathcal{F}}w_{n,m,m^{\prime}}\left(f_{n}\right),\forall n\in\mathcal{N},m,m^{\prime}\in\mathcal{M},$ (12)
@@ V-B Resource Allocation Subproblem
$\ \ \text{s.t.}\ \mathbb{P}\left\{t_{n,m_{n}}\leq D_{n}\right\}\geq 1-\varepsilon_{n},\forall n\in\mathcal{N},m_{n}\in\mathcal{G},$ (16b)
@@ V-C DNN Model Partitioning Subproblem
$\ \ \text{s.t.}\ \,\mathbb{P}\left\{\sum_{m\in\mathcal{M}}x_{n,m}t_{n,m}\leq D_{n}\right\}\geq 1-\varepsilon_{n},\forall n\in\mathcal{N},$ (24b)
@@
where $\rho>0$ is a penalty parameter. Then, the concave terms of the constraints (35c), (35d), and (35e) are linearized to obtain convex constraints for a minimization problem and to solve a sequence of convex problems successively. Specifically, at $i$th iteration, update $\left\{\mathbf{x}^{(i)},\mathbf{y}^{(i)}\right\}$ by solving the following approximate problem, which is parameterized by $\left\{\mathbf{x}^{(i-1)},\mathbf{y}^{(i-1)}\right\}$ obtained at (i1)(i-1)th iteration.
@@
where $\rho^{(i-1)}$ is the penalty parameter at the (i1)(i-1)th iteration. Problem (36) is a convex problem that can be solved efficiently by an IPT algorithm. The pseudo-code for solving problem (36) is presented in Algorithm 1. The computational complexity of solving problem (36) using an IPT algorithm is $\mathcal{O}\left(N^{3}M^{3}\right)$, and the number of iterations of the IPT algorithm is $\mathcal{O}(\sqrt{NM}\log(1/\xi))$, where $\xi$ is the convergence accuracy. Therefore, the total computational complexity of Algorithm 1 is $\mathcal{O}\left((NM)^{3.5}\log(1/\xi)\right)$ [39]. Note that the sequence solution $\left\{\mathbf{x}^{(i)}\right\}_{i=1}^{\infty}$ to problem (36) can converge to a stationary point of problem (33), as shown in [40]. Since problem (33) and problem (24) are equivalent, Algorithm 1 can also converge to a stationary point of problem (24).
@@ VI-A Simulation Setup
ViT-B/32 Jetson Nano GPU $f\in[0.1,0.8]\mathrm{GHz}$ GeForce RTX 4080
TABLE III: The parameters of ViT-B/32 on Jetson Nano GPU.
Parameter point 0 point 1 point 2 point 3 point 4 point 5 point 6
$d_{n,m}$ (MB) 0.574 0.146 0.146 0.146 0.146 0.146 0.001
$w_{n,m}$ (GFLOPs) 3.0954 3.8114 5.2435 7.3916 8.1077 8.8253
$g_{n,m}$ (FLOPs/cycle) 171.967 174.837 175.369 181.168 178.191 135.983
$v_{n,m}^{\mathrm{loc}}$ $(\mathrm{ms})^{2}$ 11.059 18.931 33.337 65.814 75.867 153.434
TABLE IV: The parameters of AlexNet on Jetson Xavier NX CPU.
@@
TABLE V: The parameters of ResNet152 on Jetson Xavier NX GPU.
@@
Three widely-used DNNs, AlexNet [23], ResNet152 [24] and ViT-B/32 [50], are considered. The three DNNs are fully deployed on mobile devices and the MEC server. The task of the mobile device is image recognition, which is extracted from the object recognition dataset CIFAR-10 [42]. The processing unit of mobile devices adopts Jetson Xavier NX CPU and GPU [43], as well as Jetson Nano GPU [51]. We assume that AlexNet is deployed on the Jetson Xavier NX CPU, ResNet152 is deployed on Jetson Xavier NX GPU, and ViT-B/32 is deployed on Jetson Nano GPU. The VM assigned to the mobile device uses the GeForce RTX 4080. Specific configurations are shown in Table II.
The energy efficiency coefficient $\kappa_{n}$ of Jetson Xavier NX CPU and GPU is evaluated using the power testing tool Tegrastats of NVIDIA [44]. Specifically, Jetson Xavier NX is first set to a fixed power consumption mode. Then, the power consumption of the CPU and GPU at different frequencies is measured, and finally, $\kappa_{n}$ is obtained based on the measured data. By estimation, the average $\kappa_{n}$ of the Jetson Xavier NX CPU, Jetson Xavier NX GPU and Jetson Nano GPU are $0.8\times 10^{-27}\mathrm{W}/(\text{cycle}/\mathrm{sec})^{3}$, $2.8\times 10^{-27}\mathrm{W}/(\text{cycle}/\mathrm{sec})^{3}$, and $3.2\times 10^{-27}\mathrm{W}/(\text{cycle}/\mathrm{sec})^{3}$, respectively.
AlexNet, ResNet152 and ViT-B/32 are divided into 8, 9 and 6 blocks, corresponding to 9, 10 and 7 partitioning points, respectively. The feature data size of each block can be calculated based on its output data shape. The mean inference time for each block is obtained through 500 experiments, and then the variance and covariance can also be calculated based on the mean and measured data. The specific parameters are shown in Table III, IV and V. Unless otherwise specified, the above parameters are used by default.
To evaluate the performance of Algorithm 1, we consider the following two policies as benchmark:
@@
    Random policy: the DNN partitioning point is randomly selected from the set of points where the feature data is smaller than the raw data.
@@
    Optimal policy: the DNN partitioning point is obtained using the exhaustive search method, which can find the optimal partitioning point, but its computational complexity is exponential.
To evaluate the performance of Algorithm 2, we consider the following two methods as benchmark:
  1. 1.
    Worst-case optimization: the upper bound of $t_{n,m}^{\mathrm{loc}}$ and $t_{n,m}^{\mathrm{vm}}$ obtained by the experiment is taken as the inference time, and the task deadline is not allowed to be violated.
  2. 2.
    Mean-value optimization: the inference time only adopts the mean value, without considering the uncertainty in inference time. This is a primitive method that bypasses the uncertainty, without any probabilistic guarantees.